a. What is the name of the thermal dynamic process as the gas goes from point A to point B? 4 b. What is the name of the t
Answer #1

a. The process happens at constant pressure, thus it is isobaric.

b. The process happens at constant volume, thus it is isochoric.

c. The work done from B to C is given by the area under the PV curve from B to C.

We can divide the area into 2 parts: a rectangle (V = 4 m3 to 5 m3 and P = 0 kPa to 2 kPa) and a triangle (V = 4 m3 to 5 m3 and P = 2 kPa to 5 kPa).

Area of the rectangle: l×b = (2-0 kPa)×(5-4 m3) = 2×103×1 = 2×103 J = 2kJ.

Area of the triangle: 1/2(b× h) = 1/2(5-4 m3) × ( 5-2 kPa) = 0.5×1×3×103 = 1.5×103J = 1.5kJ

Total area = 2kJ + 1.5 kJ = 3.5kJ.

Thus the work done is 3.5kJ